OpenStudy (anonymous):

Forrest Lumber uses the function S(t) = -135t + 810 to determine the salvage value, S(t), in dollars, of a table saw t years after its purchase. How long will it take the saw to depreciate completely?

OpenStudy (anonymous):

0= -135t +810 135t = 810 t=810/135 =6 years